OpenStudy (anonymous):

You can estimate your distance from a lightning storm by watching for a lightning flash and then counting in one-second intervals using the general rule that you are 1 km away for every 3 s you count. Explain why this rule works. Estimate the distance from a lightning storm if it takes 8 s (exactly) for the sound of the thunder to reach you after you see a flash of lightning. Assume that the temperature is 21oC. (The speed of light is approx. 2.99 x 108 m/s.)

OpenStudy (irishboy123):

The lightning bolt directly causes the thunder (by creating a violent increase in air temperature followed by expansion and then implosion); and so the lightening and thunder happen at pretty much the same time. However, you should see the lightening before you hear the thunder as light travels at 3x10^8m/s whereas sound travels at about 330m/s (ie about 1/3 of 1 km per second) --unless you are very very close to the lightening!!! You can assume that if you are a decent distance away that you see the lightening as soon as it occurs, and that thunder is on its way -- but you will have to wait a little to hear the thunder (ie experience its pressure wave) as the sound wave "only" travels at 330m/s. if the delay is 3 seconds, the thunder wve has travelled, and the lightening occurred, about 3 x 330m = 1 km away. if the delay is 10 seconds, the lightening occurred 10 x 330m = 3.3km away, and so on. [the speed of sound may be closer to 340m/s but the rounded numbers all support the same basic premise.]

OpenStudy (shamim):

Sound velocity at 0 degree celcius is 333m/s

OpenStudy (shamim):

For 1 degree celcius temperature increase velocity of sound will b increase 0.61 m/ s

OpenStudy (shamim):

So for 21 degree celcius temperature increase , velocity of sound will b increase 21*0.61=?

OpenStudy (shamim):

So velocity of sound at 21 degree celcius temperature=333+0.61*21=?

OpenStudy (shamim):

Distance travelled=velocity*time. Right?

OpenStudy (radar):

I can't add very much more than that given to you by shamim and IrishBoy123 as both have given you enough, I am saying more than likely you don't need to concern yourself with the speed of light, as seeing the lighting and the time it occurred was practically instantaneous.